OpenStudy (anonymous):

Can anyone help me with this one? I'm not quite sure of it but I think that its A... Choose the correct slope of the line that passes through the points (1, -3) and (3, -5). A: -1 B: 1 C: 0 D: 1/2

OpenStudy (anonymous):

Yes. it would be A

OpenStudy (anonymous):

Thanks again.. I just wanted to make sure before I submit this assignment...

jimthompson5910 (jim_thompson5910):

Correct, m = (y2-y1)/(x2-x1) m = (-5-(-3))/(3-1) m = (-5+3)/(3-1) m = -1

OpenStudy (anonymous):

I got another one if you could help me with it... its my last one and I cant quite seem to be able to wrap my mind around it... Choose the point-slope form of the equation below that represents the line that passes through the points (-3, 2) and (2, 1). A: y + 3 = -5(x - 2) B: y - 2 = -5(x + 3) C: y + 3 = -one fifth(x - 2) D: y - 2 = -one fifth(x + 3)

jimthompson5910 (jim_thompson5910):

what's the slope

OpenStudy (anonymous):

No slope is listed

jimthompson5910 (jim_thompson5910):

you have to find it first

OpenStudy (anonymous):

But I believe it would be 3 and 1/2...

OpenStudy (anonymous):

with a y-intercept of 21... Does that sound Correct?

jimthompson5910 (jim_thompson5910):

the slope isn't 3 1/2

OpenStudy (anonymous):

Oh snap I was looking at the wrong problem... The slope of this one should be -1/5 correct?

jimthompson5910 (jim_thompson5910):

use the slope formula

jagr2713 (jagr2713):

use slope forumla

OpenStudy (anonymous):

I think the answer is D but Im not sure right now...

jagr2713 (jagr2713):

No its A

OpenStudy (anonymous):

A? how so?

jagr2713 (jagr2713):

-3-(-3)/3-1=-2/2=-1

jagr2713 (jagr2713):

opppss sorry i am doing a wrong question lol

OpenStudy (anonymous):

Oh... Now I see what I did...

jagr2713 (jagr2713):

wait hold on let me do it

OpenStudy (anonymous):

Ok

jagr2713 (jagr2713):

Yea i got D to

OpenStudy (anonymous):

Well then I think D is my best bet... IF its wrong I got 3 more attempts so its ok if I miss 1... but thanks for helping again..

OpenStudy (anonymous):

90% and D was correct so thanks for reassuring me

jagr2713 (jagr2713):

yw